### v1

# Crystalline Spectral Form Factors

Abstract
We investigate crystalline-like behavior of the spectral form factor (SFF) in unitary quantum systems with extremely strong eigenvalue repulsion. Using a low-temperature Coulomb gas as a model of repulsive eigenvalues, we derive the Debye-Waller factor suppressing periodic oscillations of the SFF and estimate the order of its singularities at multiples of the Heisenberg time. We also reproduce this crystalline-like behavior using perturbed permutation circuits and random matrix ensembles associated with Lax matrices. Our results lay a foundation for future studies of quantum systems that exhibit intermediate level statistics between standard random matrix ensembles and permutation circuits.
Introduction— Spectral statistics and random matrix theory (RMT) are cornerstones of quantum chaos and quantum statistical mechanics [Porter, Mehta, Forrester, Haake, Stockmann, Guhr:1997ve, DAlessio:2015]. Most quantum systems fall into one of the four major RMT classes  [^1] A complete classification of random matrix ensembles reveals additional spectral features [Altland:1997, Kawabata:2018gjv]. distinguished by the probability density function of eigenvalue (energy level) spacings $s$, $P(s)\sim s^{\beta}$ as $s\to 0$. Quantum chaos is associated with Wigner-Dyson statistics, which corresponds to $\beta=1,2$ or $4$ and implies level repulsion [Bohigas:1984]. On the other hand, integrable systems lack level correlations and follow Poisson statistics with $\beta=0$ [Berry:1977a]. Since the inception of RMT in the framework of nuclear physics [Weidenmuller:2008vb], these four universal classes have been observed in a wide range of quantum systems and provided essential tools for quantum many-body dynamics [Srednicki:1994, Rigol:2008, Sachdev, Nandkishore:2014, Abanin:2018], quantum information theory [Collins:2016], quantum optics [Carmichael], quantum transport [Beenakker:1997], quantum gravity [Brezin:1978, Cotler:2016fpe, Saad:2019lba, Cotler:2020ugk, Belin:2023efa, Jafferis:2025vyp], and many other areas of physics.
However, some quantum systems fall outside the standard RMT classes with $\beta=0,1,2$ or $4$. One class of remarkable examples of non-conventional yet universal level statistics are provided by quantum cellular automata [Farrelly:2019zds, Arrighi:2019uor, Gopalakrishnan:2018, Gopalakrishnan:2018rfu, Gopalakrishnan:2018wyg, Alba:2019okd, Iaconis:2019hab, Iaconis:2021, Feldmeier:2020xxb, Hillberry:2024xpb] and random permutation circuits [Bertini:2024tuo, Szasz-Schagrin:2025dil, Bertini:2025vvz]. In a properly chosen basis, discrete unitary time evolution of these models breaks down into large periodic cycles, so their quasi-energy spectrum reduces to sets of roots of unity  [^2] Possibly shifted by a constant phase.. Once the length of the largest cycle is comparable to the dimension of the Hilbert space (or dynamics is restricted to one of the invariant subspaces), quasi-energy levels become equally spaced on the unit circle. The probability density function of such a “picket fence” ensemble is exactly zero in the vicinity of level spacing $s=0$, so this behavior corresponds to $\beta=\infty$. At the same time, these models are easily engineered and exhibit non-trivial dynamics. So, it becomes increasingly interesting to understand the emergence of quantum chaos in the entire range $\beta\in(0,\infty)$, and to find analytically tractable models that show behavior that interpolates between $\beta=\infty$ and standard RMT classes.
In this Letter, we examine three models covering the lightly explored regime between $\beta\sim 1$ and $\beta=\infty$: circular $\beta$-ensemble (Coulomb crystal of eigenvalues), perturbed permutation circuit, and random Lax matrix ensemble. We study correlations between eigenvalues using the Fourier transform of the two-level distribution, also known as the spectral form factor (SFF):
$K(t)=\left\langle\left|\mathrm{Tr}\,U(t)\right|^{2}\right\rangle=\sum_{m,n=0}^{d-1}\left\langle e^{i(E_{m}-E_{n})t}\right\rangle.$ (1)
Here, $U(t)=(U(1))^{t}$ is the evolution operator for integer time $t$, $E_{n}$ are the corresponding (quasi)energies, $d$ is the dimension of the Hilbert space, and $\langle\ldots\rangle$ is the ensemble average over statistically similar systems. Since the models that we study exhibit long-range (in energy) crystalline-like correlations of the eigenvalues of $U=U(1)$, we will employ the analogy with Bragg diffraction and treat the SFF as the structure factor of an eigenvalue crystal [Ashcroft] which is subject to fluctuations. Following this analogy, quasi-energies $E_{m}$ correspond to positions of atoms, and time $t$ corresponds to the momentum change of the scattered wave.
Figure 1: The SFF of the circular $\beta$-ensemble, which describes a Coulomb crystal of $d=512$ repulsive eigenvalues at inverse temperature $\beta=500$. We average over $10^{6}$ samples. Dashed line shows the damping Debye-Waller factor. Inset: deviation of the SFF from the plateau (zoom in the same graph).
The SFF is the simplest nontrivial quantity that captures eigenvalue correlations over the entire spectrum. In integrable systems the SFF decreases, in many cases monotonically, from $K(0)=d^{2}$ to $K(t)\to d$ as $t\to\infty$. In conventionally quantum chaotic systems, it exhibits a telltale “dip—ramp—plateau” structure, which reflects level repulsion [Haake, Stockmann, Guhr:1997ve]. However, none of these standard SFF types show structure at long times $t$ that are large multiples of $d$. The models that we study in this Letter all can exhibit crystalline-like oscillations of the SFF extending to $t\gg d$ due to extremely strong level repulsion (Fig. 1). Similarly to Bragg diffraction from a real crystal, these oscillations are damped due to the weak fluctuations in the level patterns once we move away the exact $\beta=\infty$ limit [Ashcroft]. We calculate the damping Debye-Waller factor for all mentioned models and confirm it with numerical results. We also extend this calculation to higher-order derivatives in the case of the circular $\beta$-ensemble and show that well-known features of the SFF for the standard cases $\beta=1,2,4$ are remnants of Bragg peaks.
Coulomb crystal of eigenvalues— The circular $\beta$-ensemble extends the Wigner-Dyson level statistics of unitary operators from $\beta=1,2,4$ to an arbitrary real non-negative $\beta$. The joint probability distribution for the quasi-energies $E_{n}$ of a unitary matrix drawn from the circular $\beta$-ensemble is described by the $d$-particle Coulomb gas confined to a unit circle at temperature $T=1/\beta$  [^3] The equilibration of the Coulomb gas is described by the Dyson Brownian motion [Porter, Mehta, Forrester] and reproduces probability distribution (2). We emphasize that this description does not rely on the specific properties of $\beta=1,2,4$.:
$P(E_{0},\ldots,E_{d-1})$ $=Z_{d,\beta}^{-1}\,e^{-\beta H(E_{0},\ldots,E_{d-1})},$ (2)
$H(E_{0},\ldots,E_{d-1})$ $=-\sum_{m<n}\log\left|e^{iE_{m}}-e^{iE_{n}}\right|,$
where $Z_{d,\beta}$ is the normalizing factor (partition function).
We emphasize that, in general, the probability distribution of the random matrices that realize the circular $\beta$-ensemble is not invariant under basis transformations. Only the standard values $\beta=1,2,4$ correspond to basis-invariant ensembles  [^4] Physically sound basis transformations are described by an associative division algebra over real numbers. According to the Frobenius theorem, all such algebras are isomorphic to either real, complex, or quaternion numbers. These algebras correspond to the orthogonal ($\beta=1$), unitary ($\beta=2$), and symplectic ($\beta=4$) basis transformations, respectively.. Nonetheless, level statistics (2) with arbitrary $\beta>0$ does emerge in basis-specific settings, including a class of random-walk-like models [Killip:2004, Killip:2006, CMV]. This relationship allows us to sample the circular $\beta$-ensemble at low computational cost and study its spectral properties beyond $\beta=1,2,4$. Furthermore, it opens a way for experimental tests and many-body generalizations of our models: see Supplemental Material and Ref. [Kolganov:2024nzi] for examples of quantum circuits implementing the circular $\beta$-ensemble.
To introduce the random-walk-like model of the circular $\beta$-ensemble, consider a quantum particle on a discrete interval with coordinates $n=0,1,\ldots,l-1$ and an additional internal degree of freedom with basis states $|\pm\rangle$ (thus, $d=2l$). The discrete time evolution of the particle consists of alternating shift and scattering steps:
$U=SLS^{\dagger}M\penalty 10000\ .$ (3)
The shift operator $S$ is conditional: it always flips the particle but moves it only if it occupies the $|+\rangle$ state:
$S|n\rangle|-\rangle=S|n\rangle|+\rangle\penalty 10000\ ,\quad S|n\rangle|+\rangle=|n+1\rangle|-\rangle\penalty 10000\ .$ (4)
The scattering steps are described by $2l\times 2l$ block-diagonal matrices commuting with the position operator:
$L=\Theta_{1}\oplus\Theta_{3}\oplus\ldots\oplus\Theta_{2l-1}\penalty 10000\ ,\penalty 10000\ M=\Theta_{0}\oplus\Theta_{2}\oplus\ldots\oplus\Theta_{2l-2}\penalty 10000\ ,$ (5)
where each $2\times 2$ block acts on the corresponding fixed-position subspace. Blocks $n=0,\ldots,2l-2$ are determined by complex numbers $\alpha_{n}=e^{i\phi_{n}}\cos\theta_{n}$, which control the mixing of $|\pm\rangle$ states:
$\Theta_{n}=\begin{pmatrix}e^{-i\phi_{n}}\cos\theta_{n}&\sin\theta_{n}\\ \sin\theta_{n}&-e^{i\phi_{n}}\cos\theta_{n}\end{pmatrix}$. (6)
We assume that these coefficients are independently distributed with the following probability density functions:
$P(\alpha_{n})\propto\left(1-|\alpha_{n}|^{2}\right)^{(\beta(2l-1-n)/2)-1}.$ (7)
The last block is diagonal and depends on a single real parameter uniformly distributed on the interval $(-\pi,\pi)$: $\Theta_{2l-1}=e^{i\phi_{2l-1}}\oplus 1$. This specifies the model: it is proven in [Killip:2004, Killip:2006] that the eigenvalues of unitaries (3) are distributed according to the circular $\beta$-ensemble for any $d=2l$ (the model for an odd $d$ is defined similarly).
SFF of the Coulomb gas— Before deriving the SFF of the circular $\beta$-ensemble for arbitrary $\beta$, let us first consider the low-temperature limit $\beta\to\infty$. In this limit, repulsion between Coulomb gas particles overcomes thermal fluctuations, but periodic boundary conditions keep particles confined. Hence, the gas condenses into a one-dimensional Coulomb crystal $E_{n}=(2\pi n+\phi_{d-1})/d$, where $\phi_{d-1}$ is the overall phase (center of mass position). In this limit, the SFF becomes purely periodic in $t$ with period $d$: For $t=nd$ for $n$ any integer $K(t)=d^{2}$. For all other (integer) values of $t$, $K(t)=0$. In terms of the random-walk-like model (3), the limit $\beta\to\infty$ sets all coefficients $\alpha_{n}$ to zero and turns scattering blocks into swap operators, $\Theta_{n}\to\sigma_{x}$. Hence, the evolution of the particle reduces to a cyclic permutation of basis states.
For a large but finite $\beta$, periodic oscillations of the SFF are suppressed at large $t$ by thermal fluctuations. Assuming that these fluctuations are relatively small  [^5] Expansion near infinite-temperature eigenvalues is self-consistent if the fluctuation amplitude is smaller than average level spacing, i.e., $\Delta x_{k}^{2}\approx 2\log d/\beta d^{2}<(2\pi/d)^{2}$. Hence, it is justified for $\beta>1$ and any numerically feasible system size ($d<10^{8}$)., we expand the Hamiltonian (2) near the saddle point, $E_{n}=E_{n}^{(\beta=\infty)}+x_{n}$ with $x_{n}\ll 2\pi/d$, neglect higher-order nonlinear terms, and switch to the momentum picture:
$H\approx\frac{1}{2}\sum_{p=1}^{d-1}p(d-p)\tilde{x}_{p}\tilde{x}_{-p}\penalty 10000\ ,$ (8)
where $\tilde{x}_{p}=\frac{1}{\sqrt{d}}\sum_{k=0}^{d-1}e^{-2\pi ikp/d}x_{k}$. Note that the center-of-mass degree of freedom ($p=0$ mode) decouples, so we integrate it out. In this approximation, we explicitly estimate the ensemble average in the SFF (1):
$K(t)$ $=\sum_{j,k}e^{2\pi i(k-j)t/d}\left\langle e^{it(x_{k}-x_{j})}\right\rangle$ (9)
$\approx d+d\sum_{k=1}^{d-1}e^{2\pi ikt/d}\left|Cd\sin(\pi k/d)\right|^{-4t^{2}/\beta d^{2}}.$
We assume $d\gg 1$ to approximate the sums involving two-point correlation functions. The numerical coefficient $C\approx 3.6$. Similarly to the zero-temperature SFF, “scattering amplitudes” constructively interfere near multiples of the Heisenberg time $t_{H}=d$. Hence, the finite-temperature SFF oscillates with period $t_{H}$. The effect of thermal vibrations is contained in the Debye-Waller factor $e^{-2W}$:
$K(t_{H}\tau)\approx d+e^{-2W}d^{2},$ (10)
which suppresses peaks of the SFF ($\tau=t/t_{H}=1,2,\ldots$):
$e^{-2W}\sim\begin{cases}d^{-4\tau^{2}/\beta},\quad&4\tau^{2}/\beta\ll 1\penalty 10000\ ,\\ \log(d)/d\penalty 10000\ ,\quad&4\tau^{2}/\beta=1\penalty 10000\ ,\\ (\beta/\tau^{2}d)(C\pi)^{-4\tau^{2}/\beta},\quad&4\tau^{2}/\beta\gg 1\penalty 10000\ .\end{cases}$ (11)
So, the peaks become negligible, and the SFF reaches the plateau only at times $t\gg t_{*}\approx t_{H}\sqrt{\beta/4}$. This new time scale is much larger than the Heisenberg time if $\beta\gg 1$.
The numerical calculations based on the random-walk-like model (3) agree well with theoretical estimates, e.g., see Figs. 1 and 2. In particular, we confirmed that full expression (9) and Debye-Waller factor (11) approximate the exact SFF with an error less than several per cent when $\beta>10$ and $e^{-2W}>0.01$.
Figure 2: Comparison of numerical and theoretical Debye-Waller factors of the circular $\beta$-ensemble. Points for inverse temperatures $1<\beta<20000$ are superposed.
Furthermore, the Gaussian approximation captures the order of the SFF singularities at multiples of the Heisenberg time. Differentiating (9) by $\tau=t/t_{H}$ in the limit $d\to\infty$, we find that the $n$-th derivative $K^{(n)}(\tau)$ at $\tau=1,2,\ldots$ is divergent for $n\geq\gamma$ (merely discontinuous if $n=\gamma$ and $n$ is odd), where
$\gamma=4\tau^{2}/\beta-1\penalty 10000\ .$ (12)
In particular, this formula recovers the exact results [Mehta] for $\tau=1$ ($\beta=1,2,4$) and $\tau=2$ ($\beta=4$). However, we emphasize that Gaussian approximation (9) works only for relatively small times $\tau\ll\beta$ (see Supplemental Material for details). This limitation explains the discrepancy between Eq. (12) and exact $\beta=1,2,4$ results at larger times, where the SFF in those specific cases has no singularities. Nevertheless, it leaves ample room for conspicuous crystalline-like behavior of the SFF when $\beta\gg 1$.
Figure 3: The SFF of toy model (13) interpolating between $\beta=2$ and $\beta=\infty$. We set $d=512$, $g=0.002$, and average over $10^{5}$ samples. The base permutation $S$ contains two cycles of length $d_{1}=481$ and $d_{2}=31$. Inset: The strength of the pinning potential $V(x)=\alpha x^{2}/2$ determined by fitting the SFF with Eq. (14). The permutation used in the inset contains a single cycle.
Perturbed permutation circuit— There are other models interpolating between the standard RMT universality classes. Besides empirical Brody [Brody:1973, Brody:1981] and Berry-Robnik [Berry:1984a] distributions, explicit examples of intermediate level statistics include banded matrices [Moshe:1994gc, Mirlin:1996, Bogomolny:2010, Evers:2008], Pandey-Mehta [Pandey:1982br, Lenz:1990, Lenz:1991, Dupuis:1991], Rosenzweig-Porter [Rosenzweig:1960, Kravtsov:2015], and other toy models [Izrailev:1990elx, Prosen:1993, Prosen:1994, Prosen:1997, Prosen:2002, Prosen:2018, Chan:2018dzt, Friedman:2019gyi, Sierant:2020]. However, these models are restricted to the interval between $\beta=0$ and $\beta=2$. Hence, they cannot produce the crystalline-like behavior of the SFF that is present at large $\beta$. We suggest a new simple model, which interpolates between $\beta=2$ and $\beta=\infty$:
$U=\exp(-igH)\,S\penalty 10000\ .$ (13)
Here, $H$ is a random $d\times d$ Hermitian matrix drawn from the Gaussian Unitary Ensemble (GUE) with probability density $\propto\exp\!\left[-(d/2)\mathrm{Tr}HH^{\dagger}\right]$, $S$ is a classical permutation of basis vectors supplemented with independent and uniformly distributed random phase rotations, $S|k\rangle=e^{i\phi_{k}}|\pi(k)\rangle$, and $0\leq g\leq\pi/2$ is a real parameter controlling the strength of this GUE perturbation to the permutation. For simplicity, we consider global matrices $H$ and $S$; however, model (13) is straightforwardly reproduced by a local shallow circuit  [^6] We discuss a straightforward local generalization of model (13) in the Supplemental Material. Besides, local permutation circuits with exponentially large cycles can be constructed using primitive polynomials over the binary field, e.g., see [Ippoliti:2025oro, Golomb].. We also assume that the permutation $\pi(k)$ is fixed but drawn at random. In this case, the typical length of its largest cycle is comparable to $d$ [Ford:2022], and the Heisenberg time $t_{H}=\mathcal{O}(d)$.
Similarly to the circular $\beta$-ensemble, we first consider the limit $g\to 0$. In this limit, eigenvalues of unitary (13) weakly fluctuate near the eigenvalues of the permutation matrix $S$, so we can use perturbation theory to evaluate the SFF. In general, permutation $\pi(k)$ contains $N$ cycles of length $d_{n}$, and each cycle produces a series of equally-spaced SFF peaks:
$K(t)\approx d+e^{-t^{2}g^{2}/d}\left[\left(\sum_{n=1}^{N}d_{n}\delta_{t\bmod d_{n},0}\right)^{2}-d\right].$ (14)
In other words, the model describes $N$ independent crystals of $d_{n}$ non-interacting particles (eigenvalues) pinned to the corresponding lattices by a harmonic potential $V(x)=\alpha x^{2}/2$ with $\alpha=d/g^{2}$, with temperature one.
For small but finite $g\ll 1/\sqrt{d}$, eigenvalue fluctuations are smaller than level spacings, so the structure of the SFF (14) and the damping Debye-Waller factors are qualitatively preserved, see Fig. 3. In this limit, the SFF reaches the plateau for $t\gg t_{*}\sim\sqrt{d\log d}/g\gg t_{H}$. In the opposite limit $d\gg 1/g^{2}$, repulsion between eigenvalues recovers the circular unitary ensemble (CUE), so the SFF approaches the standard “dip—ramp—plateau” form after the Thouless time $t_{\mathrm{Th}}\sim\sqrt{d}/g\ll t_{H}$. Thus as $dg^{2}$ is varied, this system interpolates between a perfect crystal that matches $\beta\rightarrow\infty$ and the $\beta=2$ CUE, but in between these two limits its SFF is different from the intermediate circular $\beta$-ensembles.
Random Lax matrix ensemble— The last example of the crystalline SFF is given by the ensemble of random Lax matrices $U$ of the Ruijsenaars-Schneider model [Bogomolny:2009, Bogomolny:2011], which also emerges as quantization of interval-exchange maps [Bogomolny:2004, Giraud:2004]:
$U_{mn}=\frac{1}{d}e^{ip_{m}}\frac{1-e^{2\pi ig}}{1-e^{2\pi i(m-n+g)/d}}\penalty 10000\ ,$ (15)
where $p_{m}$ are independent random variables uniformly distributed between $\pi$ and $\pi$, $d$ is the dimension of the Hilbert space, and parameter $0<g<1$ controls level repulsion. In the limit $g\to 0$, unitary (15) becomes diagonal and has Poisson level statistics. In the limit $g\to 1$, it turns into a cyclic shift operator supplemented with random phase rotations: $U_{mn}=e^{ip_{m}}\delta_{m+1,n}$, so has a zero-temperature eigenvalue crystal. Thus these two limits correspond to $\beta=0$ and $\beta=\infty$, respectively. For intermediate values of $g$, the spectral statistics of $U$ is described by a gas of hard rods [Bogomolny:2011]:
$P(E_{0},\ldots,E_{d-1})\propto\prod_{j=0}^{d-1}\theta\!\left(E_{j+1}-E_{j}-\frac{2\pi}{d}g\right)\penalty 10000\ ,$ (16)
where $g$ is the scaled length of the rods, and $E_{j}$ are the positions of the rod centers. Thus this system provides an interpolation between $\beta=0$ and $\beta=\infty$; again, between these two limits this system differs from the circular $\beta$- ensembles. For large $d$, the SFF behaves as:
$\frac{K(t=\tau d)}{d}\approx 1+2\mathrm{Re}\frac{1}{e^{2\pi ig\tau}\left[1+2\pi i(1-g)\tau\right]-1}\penalty 10000\ .$ (17)
This result is exact in the limit $d\to\infty$ [Bogomolny:2011] and agrees with numerical results for $d\gg 1$, e.g., see Fig. 4. Similarly to the SFF of the Coulomb gas (9), this SFF oscillates in an earlier time regime with period $t_{H}\approx d$ when $1-g\ll 1$. Not so similarly, $K(t)/d$ and all of its time derivatives remain finite for $0<g<1$ and any $\tau=t/d>0$ in the limit $d\to\infty$. For $0<t\ll d/(1-g)$ the peaks in $K(t)/d$ are of height $\sim[d/((1-g)t)]^{2}\gg 1$ and thus strong. For later times, $K(t)/d$ continues to weakly oscillate, but now with a longer period $d/g$ and amplitude $\sim d/((1-g)t)\ll 1$.
Figure 4: SFF of the random ensemble of Lax matrices (15). We set $d=512$, $g=0.98$, and average over $10^{5}$ samples.
Conclusion— We have presented three models of “crystalline” level statistics associated with oscillations of the SFF. For all three models, we have analytically derived expressions for the Debye-Waller factor, which suppresses oscillations of the SFF, and the new time scale $t_{*}\gg t_{H}$, at which the SFF approaches the plateau. The key ingredient ensuring such a crystalline-like behavior is the strong level repulsion. Our examples indicate that this behavior occurs in some broader class of chaotic quantum systems. We also emphasize that this class is not restricted to unitary discrete-time evolution, although the examples we gave are of this type. In particular, the necessary requirement for crystalline level statistics is satisfied by the Gaussian $\beta$-ensemble at large $\beta$ [Dumitriu:2002], the weakly-disordered one-dimensional Anderson localization model [Torres-Herrera:2019], an ensemble of random Lax matrices of Calogero-Moser model [Bogomolny:2009, Bogomolny:2011], and inhomogeneous discrete-time quantum walks [Cantero:2010, Linden:2009, Ambarish:2017, Gong:2022, Kempe:2003, Venegas-Andraca:2012zkr, Aharonov:1993gdz, Ambainis:2001, Ryan:2005].
We expect that our work paves a way for further investigations in various directions. First, crystalline-like eigenvalue statistics implies that dephasing in the eigenstate basis is delayed until time $t_{*}\gg t_{H}$, so it suggests a careful study of the behavior of correlation functions, entanglement, complexity, and other essential tools of quantum dynamics, for our examples between $t_{H}$ and $t_{*}$. Second, it would be interesting to find and investigate many-body systems with crystalline-like level statistics intermediate between $\beta\sim 1$ and $\beta=\infty$. A straightforward generalization of the discrete-walk-like model (3) to the case of free non-interacting fermions, which we describe in the Supplemental Meterial, resembles other toy models of the many-body SFF [Liao:2020lac, Winer:2020mdc, Flynn:2024thf, Ikeda:2024mgb] and serves as a good starting point for this endeavor. Besides, models of intermediate level statistics between $\beta=0$ and $\beta\sim 1$ play an important role for the studies of many-body-localization transitions [Oganesyan:2006, Atas:2012, Buijsman:2019, Prakash:2020psj]. It would be thrilling to find an analog of this transition between $\beta\sim 1$ and $\beta=\infty$. Finally, recent experimental advancements [Dong:2024yaf] make it tempting to experimentally confirm the crystalline-like behavior of the SFF in our models.
Acknowledgements— The authors thank Sarang Gopalakrishnan and Amit Vikram for helpful discussions. D.A.H. was supported in part by NSF QLCI grant OMA-2120757.

### v2

@@
 Our results lay a foundation for future studies of quantum systems that exhibit intermediate level statistics between standard random matrix ensembles and permutation circuits.
-Introduction— Spectral statistics and random matrix theory (RMT) are cornerstones of quantum chaos and quantum statistical mechanics [Porter, Mehta, Forrester, Haake, Stockmann, Guhr:1997ve, DAlessio:2015].
-Most quantum systems fall into one of the four major RMT classes  [^1] A complete classification of random matrix ensembles reveals additional spectral features [Altland:1997, Kawabata:2018gjv].
+Introduction— Spectral statistics and random matrix theory (RMT) are cornerstones of quantum chaos and quantum statistical mechanics [1, 2, 3, 4, 5, 6, 7].
+Most quantum systems fall into one of the four major RMT classes  [^1] A complete classification of random matrix ensembles reveals additional spectral features [98, 99].
 distinguished by the probability density function of eigenvalue (energy level) spacings $s$, $P(s)\sim s^{\beta}$ as $s\to 0$.
-Quantum chaos is associated with Wigner-Dyson statistics, which corresponds to $\beta=1,2$ or $4$ and implies level repulsion [Bohigas:1984].
-On the other hand, integrable systems lack level correlations and follow Poisson statistics with $\beta=0$ [Berry:1977a].
-Since the inception of RMT in the framework of nuclear physics [Weidenmuller:2008vb], these four universal classes have been observed in a wide range of quantum systems and provided essential tools for quantum many-body dynamics [Srednicki:1994, Rigol:2008, Sachdev, Nandkishore:2014, Abanin:2018], quantum information theory [Collins:2016], quantum optics [Carmichael], quantum transport [Beenakker:1997], quantum gravity [Brezin:1978, Cotler:2016fpe, Saad:2019lba, Cotler:2020ugk, Belin:2023efa, Jafferis:2025vyp], and many other areas of physics.
+Quantum chaos is associated with Wigner-Dyson statistics, which corresponds to $\beta=1,2$ or $4$ and implies level repulsion [9].
+On the other hand, integrable systems and localized systems lack level correlations and follow Poisson statistics with $\beta=0$ [10].
+Since the inception of RMT in the framework of nuclear physics [11], these four universal classes have been observed in a wide range of quantum systems and provided essential tools for quantum many-body dynamics [12, 13, 14, 15, 16], quantum information theory [17], quantum optics [18], quantum transport [19], quantum gravity [20, 21, 22, 23, 24, 25], and many other areas of physics.
 However, some quantum systems fall outside the standard RMT classes with $\beta=0,1,2$ or $4$.
-One class of remarkable examples of non-conventional yet universal level statistics are provided by quantum cellular automata [Farrelly:2019zds, Arrighi:2019uor, Gopalakrishnan:2018, Gopalakrishnan:2018rfu, Gopalakrishnan:2018wyg, Alba:2019okd, Iaconis:2019hab, Iaconis:2021, Feldmeier:2020xxb, Hillberry:2024xpb] and random permutation circuits [Bertini:2024tuo, Szasz-Schagrin:2025dil, Bertini:2025vvz].
-In a properly chosen basis, discrete unitary time evolution of these models breaks down into large periodic cycles, so their quasi-energy spectrum reduces to sets of roots of unity  [^2] Possibly shifted by a constant phase..
+One class of remarkable examples of non-conventional yet universal level statistics are provided by quantum cellular automata [26, 27, 28, 29, 30, 31, 32, 33, 34, 35] and random permutation circuits [36, 37, 38].
+In a properly chosen basis, discrete-time unitary evolution of these models breaks down into large periodic cycles, so their quasi-energy spectrum reduces to sets of roots of unity  [^2] Possibly shifted by a constant phase..
 Once the length of the largest cycle is comparable to the dimension of the Hilbert space (or dynamics is restricted to one of the invariant subspaces), quasi-energy levels become equally spaced on the unit circle.
 The probability density function of such a “picket fence” ensemble is exactly zero in the vicinity of level spacing $s=0$, so this behavior corresponds to $\beta=\infty$.
-At the same time, these models are easily engineered and exhibit non-trivial dynamics.
-So, it becomes increasingly interesting to understand the emergence of quantum chaos in the entire range $\beta\in(0,\infty)$, and to find analytically tractable models that show behavior that interpolates between $\beta=\infty$ and standard RMT classes.
+At the same time, these models are readily engineered and exhibit non-trivial dynamics.
+So, it becomes increasingly interesting to understand the emergence of quantum chaos in the entire range $\beta\in(0,\infty)$, and also to find analytically tractable models that show behavior that interpolates between $\beta=\infty$ and standard RMT classes.
+Figure 1: The SFF of the circular $\beta$-ensemble, which describes a Coulomb crystal of $d$ repulsive eigenvalues at inverse temperature $\beta$.
+We illustrate this for $d=512$, $\beta=500$, averaged over $10^{6}$ samples.
+Dashed line shows the damping Debye-Waller factor.
+Inset: deviation of the SFF from the plateau (zoom in the same graph).
 In this Letter, we examine three models covering the lightly explored regime between $\beta\sim 1$ and $\beta=\infty$: circular $\beta$-ensemble (Coulomb crystal of eigenvalues), perturbed permutation circuit, and random Lax matrix ensemble.
@@
 Here, $U(t)=(U(1))^{t}$ is the evolution operator for integer time $t$, $E_{n}$ are the corresponding (quasi)energies, $d$ is the dimension of the Hilbert space, and $\langle\ldots\rangle$ is the ensemble average over statistically similar systems.
-Since the models that we study exhibit long-range (in energy) crystalline-like correlations of the eigenvalues of $U=U(1)$, we will employ the analogy with Bragg diffraction and treat the SFF as the structure factor of an eigenvalue crystal [Ashcroft] which is subject to fluctuations.
+Since the models that we study exhibit long-range (in energy) crystalline-like correlations of the eigenvalues of $U=U(1)$, we will employ the analogy with Bragg diffraction and treat the SFF as the structure factor of an eigenvalue crystal which is subject to fluctuations  [^3] Note also an analogy between the eigenvalues of a random matrix ensemble and a system of spinless fermions with long-range repulsive potential [3], which also implies oscillations of the SFF [100, 101]..
 Following this analogy, quasi-energies $E_{m}$ correspond to positions of atoms, and time $t$ corresponds to the momentum change of the scattered wave.
-Figure 1: The SFF of the circular $\beta$-ensemble, which describes a Coulomb crystal of $d=512$ repulsive eigenvalues at inverse temperature $\beta=500$.
-We average over $10^{6}$ samples.
-Dashed line shows the damping Debye-Waller factor.
-Inset: deviation of the SFF from the plateau (zoom in the same graph).
 The SFF is the simplest nontrivial quantity that captures eigenvalue correlations over the entire spectrum.
 In integrable systems the SFF decreases, in many cases monotonically, from $K(0)=d^{2}$ to $K(t)\to d$ as $t\to\infty$.
-In conventionally quantum chaotic systems, it exhibits a telltale “dip—ramp—plateau” structure, which reflects level repulsion [Haake, Stockmann, Guhr:1997ve].
+In conventionally quantum chaotic systems, it exhibits a telltale “dip—ramp—plateau” structure, which reflects level repulsion [4, 5, 6].
 However, none of these standard SFF types show structure at long times $t$ that are large multiples of $d$.
@@
 1).
-Similarly to Bragg diffraction from a real crystal, these oscillations are damped due to the weak fluctuations in the level patterns once we move away the exact $\beta=\infty$ limit [Ashcroft].
+Similarly to Bragg diffraction from a real crystal, these oscillations are damped due to the weak fluctuations in the level patterns once we move away the exact $\beta=\infty$ limit [41].
 We calculate the damping Debye-Waller factor for all mentioned models and confirm it with numerical results.
@@
 Coulomb crystal of eigenvalues— The circular $\beta$-ensemble extends the Wigner-Dyson level statistics of unitary operators from $\beta=1,2,4$ to an arbitrary real non-negative $\beta$.
-The joint probability distribution for the quasi-energies $E_{n}$ of a unitary matrix drawn from the circular $\beta$-ensemble is described by the $d$-particle Coulomb gas confined to a unit circle at temperature $T=1/\beta$  [^3] The equilibration of the Coulomb gas is described by the Dyson Brownian motion [Porter, Mehta, Forrester] and reproduces probability distribution (2).
+The joint probability distribution for the quasi-energies $E_{n}$ of a unitary matrix drawn from the circular $\beta$-ensemble is described by the $d$-particle Coulomb gas confined to a unit circle at temperature $T=1/\beta$  [^4] The equilibration of the Coulomb gas is described by the Dyson Brownian motion [1, 2, 3] and reproduces probability distribution (2).
 We emphasize that this description does not rely on the specific properties of $\beta=1,2,4$.:
@@
 These algebras correspond to the orthogonal ($\beta=1$), unitary ($\beta=2$), and symplectic ($\beta=4$) basis transformations, respectively..
-Nonetheless, level statistics (2) with arbitrary $\beta>0$ does emerge in basis-specific settings, including a class of random-walk-like models [Killip:2004, Killip:2006, CMV].
+Nonetheless, level statistics (2) with arbitrary $\beta>0$ does emerge in basis-specific settings, including a class of random-walk-like models [44, 45, 46].
 This relationship allows us to sample the circular $\beta$-ensemble at low computational cost and study its spectral properties beyond $\beta=1,2,4$.
 Furthermore, it opens a way for experimental tests and many-body generalizations of our models: see Supplemental Material and Ref.
-[Kolganov:2024nzi] for examples of quantum circuits implementing the circular $\beta$-ensemble.
+[47] for examples of quantum circuits implementing the circular $\beta$-ensemble.
 To introduce the random-walk-like model of the circular $\beta$-ensemble, consider a quantum particle on a discrete interval with coordinates $n=0,1,\ldots,l-1$ and an additional internal degree of freedom with basis states $|\pm\rangle$ (thus, $d=2l$).
@@
 The last block is diagonal and depends on a single real parameter uniformly distributed on the interval $(-\pi,\pi)$: $\Theta_{2l-1}=e^{i\phi_{2l-1}}\oplus 1$.
-This specifies the model: it is proven in [Killip:2004, Killip:2006] that the eigenvalues of unitaries (3) are distributed according to the circular $\beta$-ensemble for any $d=2l$ (the model for an odd $d$ is defined similarly).
+This specifies the model: it is proven in [44, 45] that the eigenvalues of unitaries (3) are distributed according to the circular $\beta$-ensemble for any $d=2l$ (the model for an odd $d$ is defined similarly).
 SFF of the Coulomb gas— Before deriving the SFF of the circular $\beta$-ensemble for arbitrary $\beta$, let us first consider the low-temperature limit $\beta\to\infty$.
@@
 $\gamma=4\tau^{2}/\beta-1\penalty 10000\ .$ (12)
-In particular, this formula recovers the exact results [Mehta] for $\tau=1$ ($\beta=1,2,4$) and $\tau=2$ ($\beta=4$).
+In particular, this formula recovers the exact results [2, 3] for $\tau=1$ ($\beta=1,2,4$) and $\tau=2$ ($\beta=4$).
 However, we emphasize that Gaussian approximation (9) works only for relatively small times $\tau\ll\beta$ (see Supplemental Material for details).
+So, in general, the estimate (12) is guaranteed to be valid only for $\tau\ll\beta$, even though properties of a given $\beta$ can significantly extend this region  [^7] For a rational $\beta/2=p/q$ with $p$ and $q$ mutually prime, the asymptotic expansion of the density-density correlation function of the circular $\beta$-ensemble, Eq.
+(13.226) in [3], implies that the SFF has no singularities for $\tau>p$, whereas Eq.
+(12) is valid for $\tau\leq p$..
 This limitation explains the discrepancy between Eq.
 (12) and exact $\beta=1,2,4$ results at larger times, where the SFF in those specific cases has no singularities.
-Nevertheless, it leaves ample room for conspicuous crystalline-like behavior of the SFF when $\beta\gg 1$.
+Nevertheless, it leaves ample room for conspicuous crystalline-like behavior of the SFF for any $\beta\gg 1$.
 Figure 3: The SFF of toy model (13) interpolating between $\beta=2$ and $\beta=\infty$.
@@
 Perturbed permutation circuit— There are other models interpolating between the standard RMT universality classes.
-Besides empirical Brody [Brody:1973, Brody:1981] and Berry-Robnik [Berry:1984a] distributions, explicit examples of intermediate level statistics include banded matrices [Moshe:1994gc, Mirlin:1996, Bogomolny:2010, Evers:2008], Pandey-Mehta [Pandey:1982br, Lenz:1990, Lenz:1991, Dupuis:1991], Rosenzweig-Porter [Rosenzweig:1960, Kravtsov:2015], and other toy models [Izrailev:1990elx, Prosen:1993, Prosen:1994, Prosen:1997, Prosen:2002, Prosen:2018, Chan:2018dzt, Friedman:2019gyi, Sierant:2020].
+Besides empirical Brody [50, 51] and Berry-Robnik [52] distributions, explicit examples of intermediate level statistics include banded matrices [53, 54, 55, 56], Pandey-Mehta [57, 58, 59, 60], Rosenzweig-Porter [61, 62], and other toy models [63, 64, 65, 66, 67, 68, 69, 70, 71].
 However, these models are restricted to the interval between $\beta=0$ and $\beta=2$.
@@
 For simplicity, we consider global matrices $H$ and $S$; however, model (13) is straightforwardly reproduced by a local shallow circuit  [^8] We discuss a straightforward local generalization of model (13) in the Supplemental Material.
-Besides, local permutation circuits with exponentially large cycles can be constructed using primitive polynomials over the binary field, e.g., see [Ippoliti:2025oro, Golomb]..
+Besides, local permutation circuits with exponentially large cycles can be constructed using primitive polynomials over the binary field, e.g., see [102, 103]..
 We also assume that the permutation $\pi(k)$ is fixed but drawn at random.
-In this case, the typical length of its largest cycle is comparable to $d$ [Ford:2022], and the Heisenberg time $t_{H}=\mathcal{O}(d)$.
+In this case, the typical length of its largest cycle is comparable to $d$ [73], and the Heisenberg time $t_{H}=\mathcal{O}(d)$.
 Similarly to the circular $\beta$-ensemble, we first consider the limit $g\to 0$.
@@
 Thus as $dg^{2}$ is varied, this system interpolates between a perfect crystal that matches $\beta\rightarrow\infty$ and the $\beta=2$ CUE, but in between these two limits its SFF is different from the intermediate circular $\beta$-ensembles.
-Random Lax matrix ensemble— The last example of the crystalline SFF is given by the ensemble of random Lax matrices $U$ of the Ruijsenaars-Schneider model [Bogomolny:2009, Bogomolny:2011], which also emerges as quantization of interval-exchange maps [Bogomolny:2004, Giraud:2004]:
+Random Lax matrix ensemble— The last example of the crystalline SFF is given by the ensemble of random Lax matrices $U$ of the Ruijsenaars-Schneider model [74, 75], which also emerges as quantization of interval-exchange maps [76, 77]:
 $U_{mn}=\frac{1}{d}e^{ip_{m}}\frac{1-e^{2\pi ig}}{1-e^{2\pi i(m-n+g)/d}}\penalty 10000\ ,$ (15)
@@
 Thus these two limits correspond to $\beta=0$ and $\beta=\infty$, respectively.
-For intermediate values of $g$, the spectral statistics of $U$ is described by a gas of hard rods [Bogomolny:2011]:
+For intermediate values of $g$, the spectral statistics of $U$ is described by a gas of hard rods [75]:
 $P(E_{0},\ldots,E_{d-1})\propto\prod_{j=0}^{d-1}\theta\!\left(E_{j+1}-E_{j}-\frac{2\pi}{d}g\right)\penalty 10000\ ,$ (16)
@@
 $\frac{K(t=\tau d)}{d}\approx 1+2\mathrm{Re}\frac{1}{e^{2\pi ig\tau}\left[1+2\pi i(1-g)\tau\right]-1}\penalty 10000\ .$ (17)
-This result is exact in the limit $d\to\infty$ [Bogomolny:2011] and agrees with numerical results for $d\gg 1$, e.g., see Fig.
+This result is exact in the limit $d\to\infty$ [75] and agrees with numerical results for $d\gg 1$, e.g., see Fig.
 4.
@@
 We also emphasize that this class is not restricted to unitary discrete-time evolution, although the examples we gave are of this type.
-In particular, the necessary requirement for crystalline level statistics is satisfied by the Gaussian $\beta$-ensemble at large $\beta$ [Dumitriu:2002], the weakly-disordered one-dimensional Anderson localization model [Torres-Herrera:2019], an ensemble of random Lax matrices of Calogero-Moser model [Bogomolny:2009, Bogomolny:2011], and inhomogeneous discrete-time quantum walks [Cantero:2010, Linden:2009, Ambarish:2017, Gong:2022, Kempe:2003, Venegas-Andraca:2012zkr, Aharonov:1993gdz, Ambainis:2001, Ryan:2005].
+In particular, the necessary requirement for crystalline level statistics is satisfied by the Gaussian $\beta$-ensemble at large $\beta$ [78], the weakly-disordered one-dimensional Anderson localization model [79], an ensemble of random Lax matrices of Calogero-Moser model [74, 75], and inhomogeneous discrete-time quantum walks [80, 81, 82, 83, 84, 85, 86, 87, 88].
 We expect that our work paves a way for further investigations in various directions.
@@
 Second, it would be interesting to find and investigate many-body systems with crystalline-like level statistics intermediate between $\beta\sim 1$ and $\beta=\infty$.
-A straightforward generalization of the discrete-walk-like model (3) to the case of free non-interacting fermions, which we describe in the Supplemental Meterial, resembles other toy models of the many-body SFF [Liao:2020lac, Winer:2020mdc, Flynn:2024thf, Ikeda:2024mgb] and serves as a good starting point for this endeavor.
-Besides, models of intermediate level statistics between $\beta=0$ and $\beta\sim 1$ play an important role for the studies of many-body-localization transitions [Oganesyan:2006, Atas:2012, Buijsman:2019, Prakash:2020psj].
-It would be thrilling to find an analog of this transition between $\beta\sim 1$ and $\beta=\infty$.
-Finally, recent experimental advancements [Dong:2024yaf] make it tempting to experimentally confirm the crystalline-like behavior of the SFF in our models.
-Acknowledgements— The authors thank Sarang Gopalakrishnan and Amit Vikram for helpful discussions. D.A.H. was supported in part by NSF QLCI grant OMA-2120757.
+A straightforward generalization of the discrete-walk-like model (3) to the case of free non-interacting fermions, which we describe in the Supplemental Meterial, resembles other toy models of the many-body SFF [89, 90, 91, 92] and serves as a good starting point for this endeavor.
+Besides, models of intermediate level statistics between $\beta=0$ and $\beta\sim 1$ play an important role for the studies of many-body-localization transitions [93, 94, 95, 96].
+It would be interesting to find an analog of this transition between $\beta\sim 1$ and $\beta=\infty$.
+Finally, recent experimental advancements [97] make it tempting to experimentally confirm the crystalline-like behavior of the SFF in our models.
+Acknowledgements— The authors thank Sarang Gopalakrishnan, Amit Vikram, and Peter Forrester for helpful discussions. D.A.H. was supported in part by NSF QLCI grant OMA-2120757.